OpenStudy (rainbow_rocks03):

What is the area of a triangle with a base of 7 1/2 feet and a height of 4 2/3 feet?PLZ help.Medal for best response.

OpenStudy (anonymous):

14

OpenStudy (rainbow_rocks03):

14 1/6?

OpenStudy (thatonegirl_):

area of a triangle is the same as a square but divide it by 2 (same thing as multiplying by 1/2) because if u think about it 2 triangles make up a square. Area of a square or rectangle is A=bh so for the triangle it's just a=bh/2 or A=1/2bh so A=7.5x4.6667/2

OpenStudy (thatonegirl_):

You will most likely need a calculator